OpenStudy (anonymous):

2.-6,18,-54,... next three numbers, and rule

OpenStudy (anonymous):

*-3

OpenStudy (anonymous):

y=-3x

OpenStudy (anonymous):

so 162, -486, 1453... i did that in my head so maybe a little off

OpenStudy (anonymous):

its written as: \[a _{n}=2 \times (-3)^{n}\]

OpenStudy (anonymous):

err... no

OpenStudy (anonymous):

\[a _{n}=2 \times (-3)^{n-1}\]

OpenStudy (anonymous):

@ arman1002 is correct. From Mathematica:\[\text{Table}\left[2(-3)^{n-1},\{n,8\}\right]=\{2,-6,18,-54,162,-486,1458,-4374\} \]